Tim is three years older than his younger sister Kayla. The sum of the two ages is 9. How old is Tim? How old is Kayla?
Found 2 solutions by Edwin McCravy, addingup:
Answer by Edwin McCravy(20054)

Tim is three years older than his younger sister Kayla.
T = K+3
The sum of the two ages is 9.
T+K = 9

Substitute K+3 for T

K+3+K = 9
 2K+3 = 9
   2K = 6
    K = 3

Kayla is K, and K = 3, so Kayla is 3.

Tim is T, or T = K+3 = 3+3 = 6, So Tim is 6.

Edwin

Answer by addingup(3677)

T+K = 9 and
T = K+3 subtract K on both sides and add to above:
T+K = 9
+
T-K = 3
2T = 12
T = 6 Tim is 6 and since he is 3 years older than his sister, the sister is:
6-3 = 3
Now, the problem says their ages add up to 9:
Tim 6
Sis 3
-------
Total:9 Correct